Calculate much net Work per unit charge would it take to move a particle from a -15 V potential 

to a +10 V potential? Does the path of the motion matter?


1
Expert's answer
2022-04-04T09:05:39-0400

Now we know that

"q=1C"

"V_1=-15V\\V_2=10V"


"\u2206V=V_2-V_1=(10+15)=25V"

Work

"W=q\u2206V"

"W=1\times25=25J"